# ActivationReasoning: Logical Reasoning in Latent Activation Spaces

### v1

@@ 5 Discussion
Our results show that by grounding propositions in SAE-derived features, AR turns latent activations into a substrate for logical reasoning, enhancing LLM abilities across several dimensions. Across all four experiments, AR not only improves over its 8B/9B backbones but also surpasses much larger instruction-tuned models, proprietary undisclosed systems like GPT-4o, and even reasoning models like DeepSeek-R1-Distill-Llama-8B. First, in compositional reasoning (cf. PrOntoQA), AR enriches concept representations by composing new high-level concepts and applying deductive inference, maintaining accuracies above 93% across 1–5 hops, while all baselines degrade with complexity. Second, for reasoning over implicit or noisy cues (cf. R2C), AR remains robust when concepts are expressed indirectly through similes, where both small and large LLMs deteriorate, and vanilla SAEs collapse entirely. Third, on more realistic reasoning scenarios (cf. ProverQA), AR generalizes effectively to linguistically diverse problems, retaining strong performance even at the hard tier where all other models fall below 50%. Fourth, in abstract and context-sensitive safety tasks (cf. BeaverTails), AR recovers abstract notions and encodes constraints directly into the LLM’s latent space. Taken together, these findings demonstrate that structured reasoning in activation space provides a model-agnostic mechanism that not only improves reasoning performance but also increases transparency and control across synthetic, semi-natural, and real-world settings. Notably, AR achieves these improvements without significant overhead, while reasoning and CoT variants are both slower and less accurate (App.  B).
@@
Limitations. Although our implementation of AR relies on SAEs as a substrate to define logical propositions, the framework itself is not bound to them. Current SAEs offer a practical way to surface sparse, often interpretable features, but they are not always perfect. Features may be polysemous, overly context-dependent, or fail to capture abstract notions. In addition, pretrained SAEs allow only limited control over which concepts are discovered and how they are represented. Future work could explore alternative interpretability approaches, such as different autoencoding objectives, clustering methods, or other complementary representation-learning methods, that may yield representations better suited for logical reasoning in certain domains. Second, real-world reasoning challenges can become more diverse than the tasks highlighted in our work. In particular, large-scale datasets involving open-ended reasoning, long-context inference, or knowledge-intensive domains remain unexplored and remain to be evaluated in future work. Third, while the current setup relies on manually or semi-automatically defined rules and concept identification, advances in automated rule induction, probabilistic reasoning, and integration with external knowledge bases can substantially increase the adaptability of our framework. We view these directions as highly complementary and expect progress in these areas to directly enhance the applicability of our approach.
@@ 6 Conclusion
Future work includes enriching AR with automated rule discovery, probabilistic or fuzzy inference for uncertainty, integration with external knowledge bases, or visual inputs. Exploring alternative representation-learning methods beyond pretrained SAEs may yield features better suited for abstract or domain-specific reasoning. Finally, scaling AR to large-scale open-ended tasks with long contexts and diverse knowledge demands will be crucial to demonstrate robustness and generality.
@@ Acknowledgements
The authors acknowledge support from the hessian.AI Innovation Lab (funded by the Hessian Ministry for Digital Strategy and Innovation), the hessian.AISC Service Center (funded by the Federal Ministry of Education and Research, BMBF, grant No. 01IS22091), and the Center for European Research in Trusted AI (CERTAIN). This work has also benefited from the BMWE project "EU-SAI: Souveräne KI für Europa," 13IPC040G, and also from early stages of the Cluster of Excellence "Reasonable AI" funded by the German Research Foundation (DFG) under Germany’s Excellence Strategy— EXC-3057; funding will begin in 2026. This work was supported by the Priority Program (SPP) 2422 in the subproject "Optimization of active surface design of high-speed progressive tools using machine and deep learning algorithms" funded by the German Research Foundation (DFG). Additional funding was provided by the Aleph Alpha Collaboration Lab 1141.
@@ A.1 PrOntoQA Experimental Setup
For the Gemma experiments, the same parameters were used to build the concept dictionary (word-level units over all tokens), but with a different ordering strategy: unique-only features were retained. As above, representation search used the top $10$ activating features for $r_{c}$. Detection kept top-$k_{\text{in}}\!=\!7$ features per token and top-$k_{\text{multi}}\!=\!2$ features per concept for $\mathcal{R}_{\text{multi}}$, with an additional threshold $\tau\!=\!14$ imposed on activations before producing the activation matrix $A$. Steering used the same configuration as for Llama, with factor $\alpha\!=\!0.5$, and one SAE decoder weight row. As before, steering vectors for the True and False concepts were manually selected.
@@ A.2 Rail2Country Experimental Setup
To obtain the color detection concepts, we used a more restrictive search configuration, based on the top $5$ unique-only features per word-level unit. The features were selected on the evaluation set. This provided more focused latent representations for downstream reasoning.
@@
For the Gemma experiments, both the R2C-Mono setting and the R2C-Meta setting again used word-level units over all tokens. The concept dictionary was built in the same way as for Llama. Detection retained top-$k_{\text{in}}\!=\!50$ features per token and top-$k_{\text{multi}}\!=\!2$ features per concept for $\mathcal{R}_{\text{multi}}$. Steering was applied with factors $\alpha\!=\!0.8$ (R2C-Mono) and $\alpha\!=\!0.75$ (R2C-Meta), and the mean of top 10 concepts for a country from the SAE decoder. Similar to Llama, we selected the detection concepts based on the evaluation set.
@@ A.2.1 Rail2Country Dataset
The dataset consists of 300 train, eval, test samples each for both R2C-Mono and R2C-Meta. The train set was used for the initial concept search, which were then validated on the eval set. The eval set was also used to finetune the hyperparameters for this experiment. Finally, the test set was used to test the selected concepts and hyperparameters.
@@ A.3 ProverQA Experimental Setup
Since the training dataset did not contain all the required concepts, we first extracted all propositions from the dataset. These propositions were then used to locate corresponding activations in the SAE, such that each proposition provided at least one sample for identifying proposition activations. We also added 10 random sentences to increase the number of counterexamples during the concept search. Even with the context around the propositions missing, the extraction proved to be successful.
@@ Appendix E Comparison to Instruct and CoT
In addition to the main results, we report performance of instruction-tuned and CoT-augmented variants of the backbone models (Tab.  4). While both strategies yield moderate gains over the raw base models, they remain inconsistent across tasks and degrade substantially as complexity increases. For example, Llama-8B with CoT improves on PrOntoQA 1-hop but still falls below 60% at 5 hops, whereas AR sustains accuracies above 95% across all hops. Similarly, Gemma-9B with CoT achieves strong results on the easy tasks but collapses on more complex ProverQA tiers. By contrast, AR consistently outperforms all instruct and CoT variants, highlighting that activation-level reasoning provides more robust and general improvements than prompting strategies alone.
@@ Appendix F Rail2Country: Extended Results
Table 5: Rail2Country Additional Insights. Exact-match ($\uparrow$%) results on R2C-Mono and R2C-Meta variants. Top: overall results. Middle: per-simile breakdown (detection). Bottom: per-country breakdown (reasoning). AR improves both detection and reasoning over the base model.
@@
Appendix G Ethics Statement
@@
Appendix H Reproducibility Statement

### v2

@@ 5 Discussion
Our results show that by grounding propositions in SAE-derived features, AR turns latent activations into a substrate for logical reasoning, enhancing LLM abilities across several dimensions. Across all four experiments, AR not only improves over its 8B/9B backbones but also surpasses much larger instruction-tuned models, proprietary undisclosed systems like GPT-4o, and even reasoning models like DeepSeek-R1-Distill-Llama-8B. First, in compositional reasoning (cf. PrOntoQA), AR enriches concept representations by composing new high-level concepts and applying deductive inference, maintaining accuracies above 93% across 1–5 hops, while all baselines degrade with complexity. Second, for reasoning over implicit or noisy cues (cf. R2C), AR remains robust when concepts are expressed indirectly through similes, where both small and large LLMs deteriorate, and vanilla SAEs collapse entirely. Third, on more realistic reasoning scenarios (cf. ProverQA), AR generalizes effectively to linguistically diverse problems, retaining strong performance even at the hard tier where all other models fall below 50%. Fourth, in abstract and context-sensitive safety tasks (cf. BeaverTails), AR recovers abstract notions and encodes constraints directly into the LLM’s latent space. Taken together, these findings demonstrate that structured reasoning in activation space provides a model-agnostic mechanism that not only improves reasoning performance but also increases transparency and control across synthetic, semi-natural, and real-world settings. Notably, AR achieves these improvements without significant overhead, while reasoning and CoT variants are both slower and less accurate (App.  B). AR also scales favorably to long contexts (cf. Tab.  1). At inference, AR only tracks concept activations for each token, so memory and compute grow linearly with the sequence length and remain independent of the SAE dimensionality. Thus, once concepts are extracted, rule application is purely symbolic and does not degrade with increasing token distance.
@@
Limitations. Although our implementation of AR relies on SAEs as a substrate to define logical propositions, the framework itself is not bound to them. Current SAEs offer a practical way to surface sparse, often interpretable features, but they are not always perfect. Features may be polysemous, overly context-dependent, or fail to capture abstract notions, and AR already mitigates several of these effects through thresholding and multi-feature or relational representations. Pretrained SAEs also allow only limited control over which concepts are discovered and how they are represented. Importantly, AR is not tied to SAEs: any method that extracts sparse or interpretable concept signals from raw activations could be used as a replacement, and future improvements in activation-based concept discovery would transfer directly to our framework. Future work could explore alternative interpretability approaches, such as different autoencoding objectives, clustering methods, or other complementary representation-learning methods, that may yield representations better suited for logical reasoning in certain domains. Second, real-world reasoning challenges can become more diverse than the tasks highlighted in our work. In particular, large-scale datasets involving open-ended reasoning, long-context inference, or knowledge-intensive domains remain unexplored and remain to be evaluated in future work. Third, while the current setup relies on manually or semi-automatically defined rules and concept identification, advances in automated rule induction, probabilistic reasoning, and integration with external knowledge bases can substantially increase the adaptability of our framework. Notably, because AR rule specification is decoupled from model training, rules can be proposed, refined, or revised based on observed data without retraining the LLM or SAE. For instance, LLMs could generate candidate rules directly from natural language descriptions, which could then be validated against evidence and pruned as needed. We also note that the present version of AR is purely deductive; applying forward chaining over binary propositions. Moving beyond purely deductive reasoning, AR’s solver-agnostic design also allows integration of probabilistic engines such as SLIPCOVER (Bellodi & Riguzzi, 2013) or ProbFOIL (De Raedt et al., 2015) for handling uncertainty, or inductive/abductive systems (Srinivasan, 2001; Cropper & Morel, 2021) for hypothesis formation and rule discovery from sparse data. Combined with external knowledge bases, these extensions would enable AR to learn domain-specific patterns and generalize beyond manually encoded ontologies, addressing the need for adaptability in open-ended reasoning scenarios.
Finally, AR currently fixes the representation type per concept. Since different concepts benefit from different representation types (single/multi/relational), automatic selection or hybrid combinations represent a natural extension that would improve robustness to polysemantic features and complex feature interactions.
@@ 6 Conclusion
Future work includes automatic rule discovery, integration with probabilistic or inductive reasoning modules, and automatic selection of concept representation types. Exploring representation-learning approaches beyond pretrained SAEs may yield activation spaces that are better suited for abstract or domain-specific reasoning. While the present work focuses on LLMs, AR is not limited to a specific modality, since it operates fully on concept-level activations rather than a specific architecture or type of activation. Recent progress in multimodal SAEs suggests that similar reasoning could extend to vision or VLMs (Pach et al., 2025; Fel et al., 2025; Cassano et al., 2025), where AR would allow to reason over visual attributes, e.g., enforce safety constraints in image generation, or combine cross-modal cues through rule-based composition. Finally, scaling AR to large-scale open-ended tasks with long contexts and diverse knowledge demands will be key to demonstrating robustness and generality.
@@ Acknowledgements
The authors acknowledge support from the hessian.AI Service Center (funded by the Federal Ministry of Research, Technology and Space, BMFTR, grant no. 16IS22091) and the hessian.AI Innovation Lab (funded by the Hessian Ministry for Digital Strategy and Innovation, grant no. S-DIW04/0013/003), and the Center for European Research in Trusted AI (CERTAIN). This work has also benefited from the BMWE project "EU-SAI: Souveräne KI für Europa," 13IPC040G, and also from early stages of the Cluster of Excellence "Reasonable AI" funded by the German Research Foundation (DFG) under Germany’s Excellence Strategy— EXC-3057; funding will begin in 2026. This work was supported by the Priority Program (SPP) 2422 in the subproject "Optimization of active surface design of high-speed progressive tools using machine and deep learning algorithms" funded by the German Research Foundation (DFG). Additional funding was provided by the Aleph Alpha Collaboration Lab 1141. Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them. Further, this work benefited from the ICT-48 Network of AI Research Excellence Center “TAILOR” (EU Horizon 2020, GA No 952215), the Hessian research priority program LOEWE within the project WhiteBox [GA No LOEWE/ 2/13/519/03/06.001(0010)/77], the HMWK cluster projects “Adaptive Mind” and “Third Wave of AI”, and from the NHR4CES.
@@ A.1 PrOntoQA Experimental Setup
For the Gemma experiments, the same parameters were used to build the concept dictionary (word-level units over all tokens), but with a different ordering strategy: unique-only features were retained. As above, representation search used the top $10$ activating features for $r_{c}$. Detection kept top-$k_{\text{in}}\!=\!7$ features per token and top-$k_{\text{multi}}\!=\!2$ features per concept for $\mathcal{R}_{\text{multi}}$, with an additional threshold $\tau\!=\!14$ imposed on activations before producing the activation matrix $A$. Steering used the same configuration as for Llama, with factor $\alpha\!=\!0.4$, and one SAE decoder weight row. As before, steering vectors for the True and False concepts were manually selected.
@@ A.2 Rail2Country Experimental Setup
To obtain the color detection concepts, we used a more restrictive search configuration, based on the top $5$ unique-only features per word-level unit. The features were selected on the validation set. This provided more focused latent representations for downstream reasoning.
@@
For the Gemma experiments, both the R2C-Mono setting and the R2C-Meta setting again used word-level units over all tokens. The concept dictionary was built in the same way as for Llama. Detection retained top-$k_{\text{in}}\!=\!50$ features per token and top-$k_{\text{multi}}\!=\!2$ features per concept for $\mathcal{R}_{\text{multi}}$. Steering was applied with factors $\alpha\!=\!0.8$ (R2C-Mono) and $\alpha\!=\!0.75$ (R2C-Meta), and the mean of top 10 concepts for a country from the SAE decoder. Similar to Llama, we selected the detection concepts based on the validation set.
@@ A.2.1 Rail2Country Dataset
The dataset consists of 300 train, validation, test samples each for both R2C-Mono and R2C-Meta. The train set was used for the initial concept search, which were then validated on the validation set. The validation set was also used to finetune the hyperparameters for this experiment. Finally, the test set was used once at the end to evaluate the selected concepts and hyperparameters.
@@ A.3 ProverQA Experimental Setup
We first extracted all propositions from the dataset. These propositions were then used to locate corresponding activations in the SAE, such that each proposition provided at least one sample for identifying proposition activations. We also added 10 random sentences to increase the number of counterexamples during the concept search. Even without the context surrounding the propositions, the extraction proved to be successful.
@@ Appendix E Comparison to Instruct and CoT
In addition to the main results, we report performance of instruction-tuned and CoT-augmented variants of the backbone models (Tab.  4). In addition to the greedy generations, we also applied majority vote (MV) and self-consistency (SC) sampling, where we sample $4$ times for each. Furthermore, we provided the instruction-tuned variant with identical structured rules as AR, presented as a stringified Python dictionary, denoted by "(normed)", to ensure equal information. While all strategies yield moderate gains over the raw base models, they remain inconsistent across tasks and degrade substantially as complexity increases. For example, Llama-8B with CoT improves on PrOntoQA 1-hop but still falls below 60% at 5 hops, whereas AR sustains accuracies above 95% across all hops. Similarly, Gemma-9B with CoT achieves strong results on the easy tasks but collapses on more complex ProverQA tiers. By contrast, AR consistently outperforms almost all instruct and CoT variants, highlighting that activation-level reasoning provides more robust and general improvements than prompting strategies alone. It needs to be noted that the sampling variants MV and SC require substantially more tokens and thus compute to produce an answer, e.g. for Llama3.1 8B w/ instruct + CoT + SC it took $\sim$11 million tokens whereas AR only needed $\sim$2k tokens.
@@ Appendix F Rail2Country: Extended Results
Table 5: Rail2Country Ablation and Additional Insights. Exact-match ($\uparrow$%) results on R2C-Mono and R2C-Meta variants. Top: overall results. Middle: per-simile breakdown (detection). Bottom: per-country breakdown (reasoning). AR improves both detection and reasoning over the base model.
@@
Appendix G Sensitivity of AR performance to SAE placement
Prior work shows that later transformer layers encode more semantic, more monosemantic, and more stable features, whereas earlier layers mainly capture local or syntactic patterns and therefore produce noisier and less interpretable concepts (Sawmya et al., 2025; Karvonen et al., 2025). For this reason, existing SAE work typically attaches the autoencoder to late layers (often in the last third of the network), and AR follows this convention.
Importantly, AR itself is layer-agnostic. To assess sensitivity, we now include a small layer ablation where AR is applied to the adjacent layers before and after the originally selected attachment points. We selected layers $18$ and $22$ with auto-thresholding and $\mathcal{R}_{single}$ as concept representations. For the R2C Mono dataset, we selected the steering $\alpha$ as $0.8$ and $0.7$, respectively. Across these layers, performance remains highly consistent ($93.7$ for layer $18$ and $93.3$ for layer $22$), confirming that AR functions robustly across different positions in the semantic portion of the model. Our ablation with ($\mathcal{R}_{single}$) and auto-thresholding shows stable results for the adjacent layers on the R2C dataset.
Appendix H Hyperparameter Ablations
We further conduct a hyperparameter ablation on ProntoQA with Gemma2 9B to assess how AR responds to variations in its main components, namely the size of the multi-feature representation ($\mathcal{R}_{multi}$), the number of incoming features for detection (top-$k_{in}$), the steering factor $\alpha$, and the thresholding strategy. As shown in Tab.  6, AR exhibits stable performance across a broad range of settings, with strongest results for $\mathcal{R}_{multi}$ sizes between 7 and 9, top-$k_{in}=2$, and moderate steering around $\alpha=0.4$. Among these, the steering factor shows the highest sensitivity, while automatic thresholding performs on par with manually tuned values. These trends reflect that AR inherits some variability from SAE-derived features, whose activations may fluctuate or be partly polysemantic. The introduced latent representations and hyperparameters provide flexibility to mitigate such effects, ensuring reliable behavior without reliance on fragile or finely tuned configurations. Improving the stability of SAE activations remains an open challenge in mechanistic interpretability, and advances in this area are complementary to and directly beneficial for AR.
Table 6: Hyperparameter ablations on ProntoQA with Gemma2 9B.
(a) $\mathcal{R}_{multi}$ sizes.
ProntoQA
$\mathcal{R}_{multi}$ 1-hop (%) 3-hop (%) 5-hop (%)
3 38.05 35.15 37.45
5 59.25 59.80 58.60
7 93.55 93.40 93.70
9 92.00 92.20 92.55
11 81.50 82.70 81.25
(b) top-$k_{in}$ values.
ProntoQA
top-$k_{in}$ 1-hop (%) 3-hop (%) 5-hop (%)
1 81.45 81.15 81.00
2 93.55 93.40 93.70
3 65.65 63.30 63.85
(c) Steering factor $\alpha$.
ProntoQA
$\alpha$ 1-hop (%) 3-hop (%) 5-hop (%)
0.2 32.70 33.65 33.75
0.3 75.75 73.15 75.00
0.4 93.55 93.40 93.70
0.5 78.00 77.25 76.95
0.6 34.35 33.80 34.60
(d) Manual- vs. auto-thresholding.
ProntoQA
Mode 1-hop (%) 3-hop (%) 5-hop (%)
auto 93.55 93.40 93.70
$\tau=14$ 93.45 93.45 93.50
Appendix I Ethics Statement
@@
Appendix J Reproducibility Statement